OpenStudy (anonymous):

A local hamburger shop sold a combined total of 581 hamburgers and cheeseburgers on Thursday. There were 69 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Thursday?

OpenStudy (dtan5457):

c+69=581 c+c+69=581 2c+69=581 2c=512 c=256 256 cheeseburgers